Find the electric potential of a charge q with 491 nC and distance of 200 cm


1
Expert's answer
2022-05-20T04:52:59-0400
"V=k\frac{q}{r}\\
=9*10^9*\frac{491*10^{-9}}{2.00}=2210\:\rm V"